Herasa Manufacturing purchased a new machine on Jan 1, 20X1 that was built to perform one function on its assembly line. Data pertaining to this new machine are: Cost $560,000 Residual value (salvage) $54,000 Estimated service life 8 years Production output 150,000 units Using straight-line method, what is the book value at the end of the third year of ownership?
Answer #1

Cost of machine = $560,000

Residual value = $54,000

Estimated life = 8 Year

Annual depreciation = (Cost price - Residual value)/Useful life

= (560,000 - 54,000)/8

= 506,000/8

= $63,250

Accumulated depreciation at the end of three year = Annual depreciation x 3

= 63,250 x 3

= $189,750

Book value at the end of the third year = Cost of machine - Accumulated depreciation at the end of three year

= 560,000 - 189,750

= $370,250

kindly give a positive rating if you are satisfied with the solution. Please ask if you have any query related to the question, Thanks.